# All genus open mirror symmetry for the projective line

### v1

@@ 2.3. Equivariant quantum cohomology and Frobenius structure
$(\gamma_{1}\star\gamma_{2},\gamma_{3})_{\mathbb{P}^{1},T}=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\gamma_{1},\gamma_{2},\gamma_{3}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{0,3}^{\mathbb{P}^{1},T}.$
@@ 2.4. Canonical coordinates and the transition matrix
[…]
@@ 2.5. The $\mathcal{S}$-operator
$(a,\mathcal{S}(b))_{\mathbb{P}^{1},T}=(a,b)_{\mathbb{P}^{1},T}+\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}a,\frac{b}{z-\psi}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}.$
@@
    The quantum classical basis $\{\phi_{1}(q),\phi_{2}(q)\}$.
@@
    The normalized quantum classicial basis $\{\hat{\phi}_{1}(q),\hat{\phi}_{2}(q)\}$, which is self-dual.
@@
$V_{z_{1},z_{2}}(a,b)$ $=\frac{(a,b)_{\mathbb{P}^{1},T}}{z_{1}+z_{2}}+\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\frac{a}{z_{1}-\psi_{1}},\frac{b}{z_{2}-\psi_{2}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}.$
@@
$\ (a,b)_{\mathbb{P}^{1},T}+\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}a,\frac{b}{z-\psi}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}$
@@ 2.6. Equivariant $J$-function
$J(z)=1+\sum_{\alpha\in\{1,2\}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}1,\frac{\hat{\phi}_{\alpha}}{z-\psi}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}\Big{|}_{Q=1}\hat{\phi}_{\alpha}.$
@@ 2.8. The graph sum formula for descendant Gromov-Witten potential
Suppose that $02g-2+n>$ 0. Then
${\bf u}_{1},\dots,\mathbf{u}_{n}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}=\sum_{\vec{\Gamma}\in{\bf\Gamma}_{g,n}(\mathbb{P}^{1})}\frac{\omega^{\mathbf{u}}_{A}(\vec{\Gamma})}{|\mathrm{Aut}(\vec{\Gamma})|}.⟨$
3. Equivariant open Gromov-Witten theory of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$
@@ 3.1. Moduli spaces of stable maps to ($\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}$)
Let ¯$\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$ be the moduli space of degree $\beta^{\prime}$ stable maps to $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ from type $(g,h)$ bordered Riemann surfaces with $n$ interior marked points, such that the winding numbers are given by $\mu_{i}\in\mathbb{Z}$.
@@ 3.2. Equivariant open Gromov-Witten invariants
Let $\gamma_{1},\dots,\gamma_{n}\in H^{*}_{T}(\mathbb{P}^{1};\mathbb{C})$, $\beta^{\prime}\in H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. The $T$-action on $\mathbb{P}^{1}$ induces a $T$-action on the moduli space ¯$\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$. Let $F:=\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})^{T}$ be the $T$-fixed locus. The evaluation map $\mathrm{ev}_{i}:\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})\to\mathbb{P}^{1}$ is $T$-equivariant.
For $i=1,\dots,n$, let $\mathbb{L}_{i}$ be the $i$-th tautological line bundle over ¯$\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$ formed by the cotangent line at the $i$-th marked point. Define the $i$-th descendant class $\psi_{i}$ as
@@
We choose a $T$-equivariant lift $\psi_{i}^{T}\in H^{2}_{T}(\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu});\mathbb{Q})$ of $\psi_{i}$. We define the $T$-equivariant open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$:
@@
where $[F]^{\mathrm{vir}}$ is the virtual fundamental class of $F$, and $N^{\mathrm{vir}}$ is the virtual normal bundle of $F$. Since $F$ is a compact orbifold without boundary, the above integral is well-defined.
@@ 3.3. Virtual localization formula
$1\rightarrow A^{0}_{\vec{\Gamma}}\rightarrow A_{\vec{\Gamma}}\rightarrow\mathrm{Aut}(\vec{\Gamma})\rightarrow 1$
@@
¯$\overline{\mathcal{M}}_{\vec{\Gamma}}:=\prod_{v\in V^{S}(\Gamma)}\overline{\mathcal{M}}_{g_{v},F_{v}\cup S_{v}},\quad F_{\vec{\Gamma}}:=[\overline{\mathcal{M}}_{\vec{\Gamma}}/A_{\vec{\Gamma}}].$
@@
Vice versa, given a $T$-invariant stable open map $u$, one can read off its decorated graph $\vec{\Gamma}$. Therefore, we can use the decorated graphs to describe the connected components of the $T$-fixed locus of the moduli space ¯$\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$:
(3) ¯$\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})^{T}=\bigcup_{\vec{\Gamma}\in G_{g,n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})}F_{\vec{\Gamma}}.$
@@
Figure 1. The left figure is an example of the graph $\vec{\Gamma}$in $G_{g,n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$. The right figure is an example of stable map corresponding to $\vec{\Gamma}$.
@@
    where $\Lambda^{\vee}_{g}(\mathsf{u}):=\sum_{i=0}^{g}(-1)^{i}\lambda_{i}\mathsf{u}^{g-i}$ and $\lambda_{i}$ is the $i$-th Hodge class on ¯$\overline{\mathcal{M}}_{g,n}$.
@@
¯$\int_{\overline{\mathcal{M}}_{0,1}}\frac{1}{{\bf w}-\psi}={\bf w},\quad\int_{\overline{\mathcal{M}}_{0,2}}\frac{\psi_{2}^{a}}{{\bf w}-\psi_{1}}=(-{\bf w})^{a}$ $,\quad a\in\mathbb{Z}_{\geq 0},$
¯$\int_{\overline{\mathcal{M}}_{0,2}}\frac{1}{({\bf w}_{1}-\psi_{1})({\bf w}_{2}-\psi_{2})}=\frac{1}{{\bf w}_{1}+{\bf w}_{2}}$
@@ 3.4. Open/descendant correspondence
¯$\overline{\mathcal{M}}_{g,n+h}(\mathbb{P}^{1},\beta)^{T}=\bigcup_{\hat{\Gamma}\in G_{g,n+h}(\mathbb{P}^{1},\beta)}F_{\hat{\Gamma}}.$
@@
by cutting all half-edges $h_{j}$ of $\vec{\Gamma}$, and then labeling the attached $\bullet$ vertex of $h_{j}$ with $n+j$. Because the half-edges $h_{j}$ are labeled, we have $\mathrm{Aut}(\vec{\Gamma})=\mathrm{Aut}(\hat{\Gamma})$.
@@
$\langle\tau_{a_{1}}(\gamma_{1})\dots\tau_{a_{n}}(\gamma_{n})\rangle^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}_{g,\beta^{\prime},\vec{\mu}}=\prod_{j\in J_{-}}D^{1}(-\mu_{j})\prod_{j\in J_{+}}D^{2}(\mu_{j})$
[¯$\cdot\int_{[\overline{\mathcal{M}}_{g,n+h}(\mathbb{P}^{1},\beta)]^{{\mathrm{vir}},T}}\frac{\prod_{i=1}^{n}\psi_{i}^{a_{i}}\mathrm{ev}_{i}^{*}(\gamma_{i})\prod_{j\in J_{-}}\mathrm{ev}_{n+j}^{*}\phi_{1}\prod_{j\in J_{+}}\mathrm{ev}_{n+j}^{*}\phi_{2}}{\prod_{j=1}^{h}\frac{\mathsf{v}}{\mu_{j}}(\frac{\mathsf{v}}{\mu_{j}}-\psi_{n+j})}.$
@@ 3.5. Generating function of open Gromov-Witten invariants
    $=\sum_{\alpha\in\{1,2\}}\Big{(}\frac{1}{\Delta^{\alpha}}\tilde{\xi}^{\alpha}_{-2}(X)+(\frac{t^{0}}{\Delta^{\alpha}}+\frac{t^{1}}{2})\tilde{\xi}^{\alpha}_{-1}(X)+\sum_{k\geq 0}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\phi_{\alpha}\psi^{k}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{0,1}^{\mathbb{P}^{1},T}\tilde{\xi}^{\alpha}_{k}(X)\Big{)}$
@@
    $=\sum_{k_{1},k_{2}\geq 0}\sum_{\alpha_{1},\alpha_{2}\in\{1,2\}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\phi_{\alpha_{1}}\psi^{k_{1}},\phi_{\alpha_{2}}\psi^{k_{2}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}\tilde{\xi}_{k_{1}}^{\alpha_{1}}(X_{1})\tilde{\xi}_{k_{2}}^{\alpha_{2}}(X_{2})$
    $=[z_{1}^{-1}z_{2}^{-1}]\sum_{\alpha_{1},\alpha_{2}\in\{1,2\}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\frac{\phi_{\alpha_{1}}}{z_{1}-\psi_{1}},\frac{\phi_{\alpha_{2}}}{z_{2}-\psi_{2}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}\tilde{\xi}^{\alpha_{1}}(z_{1},X_{1})\tilde{\xi}^{\alpha_{2}}(z_{2},X_{2}),$
@@
    $=\sum_{k_{1},\dots,k_{n}\geq 0}\sum_{\alpha_{1},\dots,\alpha_{n}\in\{1,2\}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\phi_{\alpha_{1}}\psi^{k_{1}},\dots,\phi_{\alpha_{n}}\psi^{k_{n}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}\prod_{j=1}^{n}\tilde{\xi}_{k_{j}}^{\alpha_{j}}(X_{j})$
    $=[z_{1}^{-1}\dots z_{n}^{-1}]\sum_{\alpha_{1},\dots,\alpha_{n}\in\{1,2\}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\frac{\phi_{\alpha_{1}}}{z_{1}-\psi_{1}},\dots,\frac{\phi_{\alpha_{n}}}{z_{n}-\psi_{n}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}\prod_{j=1}^{n}\tilde{\xi}^{\alpha_{j}}(z_{j},X_{j}).$
@@
(5) (~$(\tilde{\mathcal{L}}^{O})^{\beta}_{k}(l_{j})=[z^{k}](\sum_{\alpha,\gamma\in\{1,2\}}(\tilde{\xi}^{\alpha}(z,X_{j})S^{\hat{\underline{\gamma}}}_{\ \alpha})_{+}R(-z)_{\gamma}^{\ \beta}).$
@@
Let $\omega_{A}^{O}(\vec{\Gamma})=\widetilde{\omega}_{A}^{O}(\vec{\Gamma})\big{|}_{Q=1}$. For $02g-2+n>$ 0, we have

### v2

@@ 2.3. Equivariant quantum cohomology and Frobenius structure
$(\gamma_{1}\star\gamma_{2},\gamma_{3})_{\mathbb{P}^{1},T}=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\gamma_{1},\gamma_{2},\gamma_{3}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,3}^{\mathbb{P}^{1},T}.$
@@ 2.4. Canonical coordinates and the transition matrix
$(\Psi^{-1})_{\alpha}^{\ 0}=\frac{\sqrt{\Delta^{\alpha}(q)}}{2},\quad(\Psi^{-1})_{\alpha}^{\ 1}=\frac{1}{\sqrt{\Delta^{\alpha}(q)}}.$
@@ 2.5. The $\mathcal{S}$-operator
$(a,\mathcal{S}(b))_{\mathbb{P}^{1},T}=(a,b)_{\mathbb{P}^{1},T}+\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}a,\frac{b}{z-\psi}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}.$
@@
    The quantum canonical basis $\{\phi_{1}(q),\phi_{2}(q)\}$.
@@
    The normalized quantum canonical basis $\{\hat{\phi}_{1}(q),\hat{\phi}_{2}(q)\}$, which is self-dual.
@@
$V_{z_{1},z_{2}}(a,b)$ $=\frac{(a,b)_{\mathbb{P}^{1},T}}{z_{1}+z_{2}}+\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\frac{a}{z_{1}-\psi_{1}},\frac{b}{z_{2}-\psi_{2}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}.$
@@
$\ (a,b)_{\mathbb{P}^{1},T}+\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}a,\frac{b}{z-\psi}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}$
@@ 2.6. Equivariant $J$-function
$J(z)=1+\sum_{\alpha\in\{1,2\}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}1,\frac{\hat{\phi}_{\alpha}}{z-\psi}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}\Big|_{Q=1}\hat{\phi}_{\alpha}.$
@@ 2.8. The graph sum formula for descendant Gromov-Witten potential
Suppose that 2g2+n>02g-2+n>0. Then
$\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}{\bf u}_{1},\dots,\mathbf{u}_{n}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}=\sum_{\vec{\Gamma}\in{\bf\Gamma}_{g,n}(\mathbb{P}^{1})}\frac{\omega^{\mathbf{u}}_{A}(\vec{\Gamma})}{|\mathrm{Aut}(\vec{\Gamma})|}.$
3. Equivariant open Gromov-Witten theory of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ via integration over the fixed locus
@@ 3.1. Moduli spaces of stable maps to ($\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}$)
Let $\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$ be the moduli space of degree $\beta^{\prime}$ stable maps to $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ from type $(g,h)$ bordered Riemann surfaces with $n$ interior marked points, such that the winding numbers are given by $\mu_{i}\in\mathbb{Z}$.
@@ 3.2. Equivariant open Gromov-Witten invariants
Let $\gamma_{1},\dots,\gamma_{n}\in H^{*}_{T}(\mathbb{P}^{1};\mathbb{C})$, $\beta^{\prime}\in H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. The $T$-action on $\mathbb{P}^{1}$ induces a $T$-action on the moduli space $\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$. Let $F:=\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})^{T}$ be the $T$-fixed locus. The evaluation map $\mathrm{ev}_{i}:\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})\to\mathbb{P}^{1}$ is $T$-equivariant.
For $i=1,\dots,n$, let $\mathbb{L}_{i}$ be the $i$-th tautological line bundle over $\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$ formed by the cotangent line at the $i$-th marked point. Define the $i$-th descendant class $\psi_{i}$ as
@@
We choose a $T$-equivariant lift $\psi_{i}^{T}\in H^{2}_{T}(\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu});\mathbb{Q})$ of $\psi_{i}$. We define the $T$-equivariant open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ via integration over the fixed locus:
@@
where $F]^{\mathrm{vir}}$ is the virtual fundamental class of $F$, and $N^{\mathrm{vir}}$ is the virtual normal bundle of $F$. Since $F$ is a compact orbifold without boundary, the above integral is well-defined.
@@ 3.3. Virtual localization formula
$1\rightarrow A^{0}_{\vec{\Gamma}}\rightarrow A_{\vec{\Gamma}}\rightarrow\mathrm{Aut}(\vec{\Gamma})\rightarrow 1.$
@@
$\overline{\mathcal{M}}_{\vec{\Gamma}}:=\prod_{v\in V^{S}(\Gamma)}\overline{\mathcal{M}}_{g_{v},F_{v}\cup S_{v}},\quad F_{\vec{\Gamma}}:=[\overline{\mathcal{M}}_{\vec{\Gamma}}/A_{\vec{\Gamma}}].$
@@
Vice versa, given a $T$-invariant stable open map $u$, one can read off its decorated graph $\vec{\Gamma}$. Therefore, we can use the decorated graphs to describe the connected components of the $T$-fixed locus of the moduli space $\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$:
(3) $\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})^{T}=\bigcup_{\vec{\Gamma}\in G_{g,n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})}F_{\vec{\Gamma}}.$
@@
Figure 1. The left figure is an example of the graph $\vec{\Gamma}$ $\in$ $G_{g,n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$. The right figure is an example of stable map corresponding to $\vec{\Gamma}$.
@@
    where $\Lambda^{\vee}_{g}(\mathsf{u}):=\sum_{i=0}^{g}(-1)^{i}\lambda_{i}\mathsf{u}^{g-i}$ and $\lambda_{i}$ is the $i$-th Hodge class on $\overline{\mathcal{M}}_{g,n}$.
@@
$\int_{\overline{\mathcal{M}}_{0,1}}\frac{1}{{\bf w}-\psi}={\bf w},\quad\int_{\overline{\mathcal{M}}_{0,2}}\frac{\psi_{2}^{a}}{{\bf w}-\psi_{1}}=(-{\bf w})^{a}$ $,\quad a\in\mathbb{Z}_{\geq 0},$
$\int_{\overline{\mathcal{M}}_{0,2}}\frac{1}{({\bf w}_{1}-\psi_{1})({\bf w}_{2}-\psi_{2})}=\frac{1}{{\bf w}_{1}+{\bf w}_{2}}$ .
@@ 3.4. Open/descendant correspondence
$\overline{\mathcal{M}}_{g,n+h}(\mathbb{P}^{1},\beta)^{T}=\bigcup_{\hat{\Gamma}\in G_{g,n+h}(\mathbb{P}^{1},\beta)}F_{\hat{\Gamma}}$.
@@
[…]
@@
$\langle\tau_{a_{1}}(\gamma_{1})\dots\tau_{a_{n}}(\gamma_{n})\rangle^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}_{g,\beta^{\prime},\vec{\mu}}=\prod_{j\in J_{-}}D^{1}(-\mu_{j})\prod_{j\in J_{+}}D^{2}(\mu_{j})\ \ \ \ \ \ \ \ \ \ \ \ \$
$\cdot\int_{[\overline{\mathcal{M}}_{g,n+h}(\mathbb{P}^{1},\beta)]^{{\mathrm{vir}},T}}\frac{\prod_{i=1}^{n}\psi_{i}^{a_{i}}\mathrm{ev}_{i}^{*}(\gamma_{i})\prod_{j\in J_{-}}\mathrm{ev}_{n+j}^{*}\phi_{1}\prod_{j\in J_{+}}\mathrm{ev}_{n+j}^{*}\phi_{2}}{\prod_{j=1}^{h}\frac{\mathsf{v}}{\mu_{j}}(\frac{\mathsf{v}}{\mu_{j}}-\psi_{n+j})}.$
@@ 3.5. Generating function of open Gromov-Witten invariants
    $=\sum_{\alpha\in\{1,2\}}\Big(\frac{1}{\Delta^{\alpha}}\tilde{\xi}^{\alpha}_{-2}(X)+(\frac{t^{0}}{\Delta^{\alpha}}+\frac{t^{1}}{2})\tilde{\xi}^{\alpha}_{-1}(X)+\sum_{k\geq 0}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\phi_{\alpha}\psi^{k}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,1}^{\mathbb{P}^{1},T}\tilde{\xi}^{\alpha}_{k}(X)\Big)$
@@
    $=\sum_{k_{1},k_{2}\geq 0}\sum_{\alpha_{1},\alpha_{2}\in\{1,2\}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\phi_{\alpha_{1}}\psi^{k_{1}},\phi_{\alpha_{2}}\psi^{k_{2}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}\tilde{\xi}_{k_{1}}^{\alpha_{1}}(X_{1})\tilde{\xi}_{k_{2}}^{\alpha_{2}}(X_{2})$
    $=[z_{1}^{-1}z_{2}^{-1}]\sum_{\alpha_{1},\alpha_{2}\in\{1,2\}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\frac{\phi_{\alpha_{1}}}{z_{1}-\psi_{1}},\frac{\phi_{\alpha_{2}}}{z_{2}-\psi_{2}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}\tilde{\xi}^{\alpha_{1}}(z_{1},X_{1})\tilde{\xi}^{\alpha_{2}}(z_{2},X_{2}),$
@@
    $=\sum_{k_{1},\dots,k_{n}\geq 0}\sum_{\alpha_{1},\dots,\alpha_{n}\in\{1,2\}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\phi_{\alpha_{1}}\psi^{k_{1}},\dots,\phi_{\alpha_{n}}\psi^{k_{n}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}\prod_{j=1}^{n}\tilde{\xi}_{k_{j}}^{\alpha_{j}}(X_{j})$
    $=[z_{1}^{-1}\dots z_{n}^{-1}]\sum_{\alpha_{1},\dots,\alpha_{n}\in\{1,2\}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\frac{\phi_{\alpha_{1}}}{z_{1}-\psi_{1}},\dots,\frac{\phi_{\alpha_{n}}}{z_{n}-\psi_{n}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}\prod_{j=1}^{n}\tilde{\xi}^{\alpha_{j}}(z_{j},X_{j}).$
@@
(5) $(\tilde{\mathcal{L}}^{O})^{\beta}_{k}(l_{j})=[z^{k}](\sum_{\alpha,\gamma\in\{1,2\}}(\tilde{\xi}^{\alpha}(z,X_{j})S^{\hat{\underline{\gamma}}}_{\ \alpha})_{+}R(-z)_{\gamma}^{\ \beta}).$
@@
Let $\omega_{A}^{O}(\vec{\Gamma})=\widetilde{\omega}_{A}^{O}(\vec{\Gamma})\big|_{Q=1}$. For 2g2+n>02g-2+n>0, we have